a gas company experiences variation in both demand for parts an in lead time in ordering them, the supply office notices normally distributed daily demand rate with a mean of 43 parts, and a standard deviation of 5.2. the replenishment lead time is normally distributed with a mean of 9 days and standard deviation of 4 days.

A) if the daily demand and the lead time are independent, what reorder point should be established to provided a 95% level per order cycle?
Answer #1

Demand (d) = 43

standard deviation (s) = 5.2

Lead time (LT) = 9

standard deviation of lead time (ss) = 4

Service level is 95%. z value for 95% service level is 1.64

Reorder point = d*LT + z*sqrt(LT*(s)^2 + (d)^2*(ss)^2) = 43*9 + 1.64*sqrt(9*(5.2)^2 + (43)^2*(4)^2) = 670.2